OpenStudy (anonymous):

Need help with this equation. Just need to make sure I have the answer correct. 3xsquared-4x=2xsqaured +2 using the quadratic formula.

hartnn (hartnn):

what values did u take for a,b,c ?

OpenStudy (anonymous):

a=1, b=-4, c= -2

hartnn (hartnn):

thats correct :) what did u get ?

OpenStudy (anonymous):

4 plus or minus square root of 24 divided by 2 (but that can also be reduced)

hartnn (hartnn):

yes, tried reducing ?

OpenStudy (anonymous):

2+ square root of 24, and 2 mius square root of 24

hartnn (hartnn):

nopes. \(\huge \sqrt{24}=\sqrt{4(6)}=2\sqrt6\) so \(\huge\frac{4\pm 2\sqrt{6}}{2}=??\)

OpenStudy (anonymous):

Ok. I know how to break down square roots, so I have to do that each time. And now, would the answer be 2 plus or minus 2 square roots of 6?

hartnn (hartnn):

no! \(\huge\frac{4\pm 2\sqrt{6}}{2}=\frac{4}{2}\pm \frac{2\sqrt6}{2}=?\)

OpenStudy (anonymous):

Got ya. The divisor should apply to both the whole number and square root which means that this is now 2 plus or minus 1 on the square root of 6.

hartnn (hartnn):

yes, u have to separate denominator.example: \(\huge\frac{10+20}{10}\ne 1+20\) \(\huge\frac{10+20}{10}= \frac{10}{10}+\frac{20}{10}\)

hartnn (hartnn):

final answer here \(\huge 2\pm\sqrt6\)

OpenStudy (anonymous):

Thanks for you help!

hartnn (hartnn):

your welcome :)